According to a National Health Survey, American men’s heights are normally distributed with a mean given by 𝜇 = 69.7 inches and a standard deviation given by 𝜎 = 2.8 inches.



a). If a man is randomly selected, find the probability that his height is more 72 inches.


b). If a man is randomly selected, find the probability that his height is between 68 and 72 inches.


1
Expert's answer
2022-05-05T12:43:09-0400

a)


"P(X>72)=1-P(Z\le\dfrac{72-69.7}{2.8})\approx0.2057"

b)


"P(68<X<72)=P(Z<\dfrac{72-69.7}{2.8})"

"-P(Z\le\dfrac{68-69.7}{2.8})\approx0.5224"